# Adversarial Generation and Collaborative Evolution of Safety-Critical Scenarios for Autonomous Vehicles

### v1

@@ 4.4 Generalization Ability Analysis
Model Generalization. Beyond RL-based AV models, we further evaluate our generated scenarios on LMDrive [41], a large vision-language model for AV deployed on the CARLA Leaderboard [45]. LMDrive navigates by following natural language instructions sequentially, using the multi-view camera and Lidar perception for scene understanding and planning. To accommodate its instruction-driven execution mode, we redesign the test routes into multi-instruction sequences. Evaluation follows LMDrive’s original metrics: Route Completion (RC), Infraction Score (IS), and Driving Score (DS). We evaluate LMDrive under three increasingly challenging settings: ❶ ego-only benign routes as a baseline, ❷ meta-scenarios with a single adversarial agent, and ❸ full adversarial scenarios generated by our framework, including the perturbed background vehicle. As shown in Tab. 3, LMDrive’s performance drops from 87.7 DS in the benign case to 83.7 in meta-scenarios and further to 80.4 under full adversarial conditions. These results demonstrate that our generated scenarios significantly stress LMDrive’s planning capability, especially under rare or occluded interactions.
@@
Simulator Generalization. We validated the generalizability of ScenGE in MetaDrive [18] simulator, a lightweight platform for RL research. Despite substantial differences in rendering and physics, ScenGE’s modular design facilitated a successful port. As shown in Fig. 4, our framework successfully generated effective safety-critical scenarios that induced high CR in MetaDrive. This result demonstrates the broad applicability of our approach.
4.5 Training on the Generated Scenarios
@@
Robustness evaluation on real-world data. We further evaluated our adversarially trained model on real-world data from the nuScenes dataset [1] to address the visual domain gap between simulation and reality. In particular, we first manually select 140 scenario segments (2700 images) with latent risks, such as a pedestrian standing by the roadside; subsequently, we evaluate both the original RL model and the enhanced model on these real-world image sequences. The enhanced model achieved a 21.7% lower Euclidean distance to the ground truth trajectory than the original model, indicating safer and more stable decisions. This result demonstrates that the robustness gained from our scenarios transfers effectively to real-world perception data.

### v2

@@ 4.4 Generalization Ability Analysis
Model Generalization. Beyond RL-based AV models, we further evaluate our generated scenarios on LMDrive shao2024lmdrive , a large vision-language model for AV deployed on the CARLA Leaderboard leaderboard . LMDrive navigates by following natural language instructions sequentially, using the multi-view camera and Lidar perception for scene understanding and planning. To accommodate its instruction-driven execution mode, we redesign the test routes into multi-instruction sequences. Evaluation follows LMDrive’s original metrics: Route Completion (RC), Infraction Score (IS), and Driving Score (DS). We evaluate LMDrive under three increasingly challenging settings: ❶ ego-only benign routes as a baseline, ❷ meta-scenarios with a single adversarial agent, and ❸ full adversarial scenarios generated by our framework, including the perturbed background vehicle. As shown in Tab. 3, LMDrive’s performance drops from 87.7 DS in the benign case to 83.7 in meta-scenarios and further to 80.4 under full adversarial conditions. These results demonstrate that our generated scenarios significantly stress LMDrive’s planning capability, especially under rare or occluded interactions.
Simulator Generalization. We validated the generalizability of ScenGE in MetaDrive li2021metadrive simulator, a lightweight platform for RL research. Despite substantial differences in rendering and physics, ScenGE’s modular design facilitated a successful port. As shown in Fig. 4, our framework successfully generated effective safety-critical scenarios that induced high CR in MetaDrive. This result demonstrates the broad applicability of our approach.
4.5 Training on the Generated Scenarios
@@
Robustness evaluation on real-world data. We further evaluated our adversarially trained model on real-world data from the nuScenes dataset caesar2020nuscenes to address the visual domain gap between simulation and reality. In particular, we first manually select 140 scenario segments (2700 images) with latent risks, such as a pedestrian standing by the roadside; subsequently, we evaluate both the original RL model and the enhanced model on these real-world image sequences. The enhanced model achieved a 21.7% lower Euclidean distance to the ground truth trajectory than the original model, indicating safer and more stable decisions. This result demonstrates that the robustness gained from our scenarios transfers effectively to real-world perception data.